# PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games

### v1

@@ 3.1 Experimental Setup
We conducted experiments with GPT-3.5 for conversation and the GPT Embedding Model for memory retrieval via the Azure API in Jan-Mar 2024, using the default model versions gpt-35-turbo-16k 0613 and text-embedding-ada-002. To minimise randomness, we conducted the evaluation experiments 3 times and report the average and variance. The detailed introduction to the experiment settings is in Appendix C.1.
@@
For baselines, we compare our approach with other multi-agent algorithms designed for multiplayer deduction games, such as Werewolf (Xu et al., 2023a), Objective-Guided Chain of Thought (O-CoT) (Park et al., 2023; Zhao et al., 2024b), and ThinkThrice (Wu et al., 2024). Additionally, we assess the performance of agents that do not actively participate in the game but have direct access to either their own script (Single Script Access) or the scripts of all agents (Full Script Access). In werewolf, questions are chosen from a role-specific predefined list to facilitate game progression, alongside questions generated based on the current scenario. In O-CoT, agent interactions are driven by their set objectives. ThinkThrice has its agents crafting questions from retrieved memory and the current scenario. However, the efficiency of exploration with the questioning strategies is low, as evidenced by their performance, which falls significantly short of full script access. In contrast, our method leverages sensor data to identify the optimal domain to question and guide the generation of questions. The detailed comparison is in Appendix C.2.
WellPlay Dataset
We create an evaluation dataset, called the WellPlay Dataset, based on the existing dataset Conan (Zhao et al., 2024c), originated from background narratives created for MMGs, which provides annotated relationships between characters. We aim to concentrate on the most challenging and significant relationships that assess the agents’ comprehension of the scenario. Our approach is grounded in the original game’s objectives and informed by a trial round involving human players. This method enabled us to formulate the following types of evaluation questions in line with the human players:
1. Objective. Including shared objectives, such as identifying the perpetrator(s), and individual objectives, such as determining who stoles wallet, for each character in the game.
2. Reasoning. This entails questions that delve into the reasoning behind provided answers, relating to agents’ objectives, including: Who; What (the nature of the incident, such as murder, theft, disappearance, explosion); When (the time of the incident); Where(the location of the incident); Cause (e.g., shooting, poisoning, stabbing); Motive (e.g., crime of passion, vendetta, financial conflicts, manslaughter).
3. Relations. This includes interpersonal relationships between victims and other characters, as well as relationships among suspects.
To establish a quantifiable evaluation method, we employ multiple-choice questions focusing on factual information. This dataset encompasses 12 MMGs, comprising a total of 1,482 evaluation questions. On average, each game features 5.67 agents and 1.75 victims. More detailed statistics are available in the appendix B.1.
Evaluation Metrics
We adopted the scoring system used in the game to evaluate agents’ performance: Awarding 10 points for achieving an objective, 5 points for correct reasoning, and 2 points for providing additional information based on the players’ understanding of the current situation. However, instead of using open-ended questions with human players, we assess each agent’s performance using our constructed dataset, WellPlay, to facilitate replication and comparison for future research. WellPlay comprises three types of questions: objective, reasoning, and relations, which are assigned 10, 5, and 2 points, respectively. We calculate the agent’s final score as Awarded PointsTotal $\frac{\text{Awarded Points}}{\text{Total Points}}$ to determine its overall performance. Additionally, we assess the accuracy of each question type. For the computation of the overall score for performance assessment, we have utilized weighted mean and weighted standard deviation methodologies. The detailed algorithm for these calculations is in Appendix B.2.
@@ 3.2 Results
Table 1 presents the evaluation of agents’ performance across 12 unique games of varying complexity and settings. The number of evaluation questions (“#QA”) varies based on script complexity, with more complex scripts generating a larger volume of questions. The SSA (Single Script Access) measures agents’ performance with access to only their own script. This setup is designed to represent the starting point for searching. The FSA (Full Script Access) measures performance when agents have access to all agents’ scripts, representing the ideal search endpoint. Ideally, FSA should aim to achieve an accuracy of 1. However, in practice, its effectiveness is often limited by the deductive capabilities of the underlying base model. Despite this, FSA still scores significantly higher than SSA. In addition, FSA also serves as an indicator of the complexity of the script.
We can observe that PLAYER* exhibits superior performance other baselinse across all evaluation questions, demonstrating its enhanced understanding of the search space through interactions with other agents. Additionally, PLAYER* significantly outperforms others in objective questions, even surpassing FSA, which has access to all information. This showcases the effectiveness of PLAYER* in refining information and dynamically narrowing down the search domain to achieve the target objective. Despite some baselines exhibiting competitive performance in reasoning or relation questions, their significant drop in performance concerning objectives indicates a critical limitation: an inability to effectively utilise the collected information to reach correct conclusions or achieve game-specific goals.
We also evaluate the performance of our framework on English scripts in Appendix C.3 and open-source LLMs in Appendix C.4. The detailed dialogue history and evaluation records are available in the GitHub link provided previously.
Script Evaluation #QA SSA FSA Agent’s Response After Playing the Game
Werewolf O-CoT ThinkThrice PLAYER*
Death Wears White (9 players, 1 victim) Objective $10$$233_{\pm.058}.$$467_{\pm.116}.$$333_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.116}.$
Ghost Revenge (7 players, 3 victims) Objective $19$$333_{\pm.030}.$$334_{\pm.132}.$$368_{\pm.052}.$$316_{\pm.190}.$$404_{\pm.080}.$$509_{\pm.133}.$
Danshui Villa (7 players, 2 victims) Objective $12$$389_{\pm.048}.$$222_{\pm.127}.$$167_{\pm.000}.$$305_{\pm.048}.$$306_{\pm.096}.$$361_{\pm.127}.$
Unfinished Love (7 players, 2 victims) Objective $12$$167_{\pm.000}.$$167_{\pm.083}.$$167_{\pm.144}.$$278_{\pm.048}.$$111_{\pm.127}.$$222_{\pm.127}.$
Cruise Incident (5 players, 1 victim) Objective $4$$167_{\pm.289}.$$250_{\pm.250}.$$500_{\pm.250}.$$500_{\pm.000}.$$583_{\pm.382}.$$667_{\pm.144}.$
Sin (4 players, 1 victim) Objective $3$$000_{\pm.000}.$$333_{\pm.334}.$$111_{\pm.192}.$$111_{\pm.192}.$$111_{\pm.192}.$$333_{\pm.334}.$
Deadly Fountain (4 players, 1 victim) Objective $3$$000_{\pm.000}.$$556_{\pm.193}.$$000_{\pm.000}.$$111_{\pm.192}.$$000_{\pm.000}.$$111_{\pm.192}.$
Unbelievable Incident (5 players, 1 victim) Objective $4$$083_{\pm.144}.$$083_{\pm.144}.$$333_{\pm.144}.$$167_{\pm.144}.$$083_{\pm.144}.$$250_{\pm.000}.$
Desperate Sunshine (4 players, 1 victim) Objective $3$$333_{\pm.000}.$$556_{\pm.193}.$$556_{\pm.193}.$$333_{\pm.000}.$$667_{\pm.334}.$$778_{\pm.192}.$
Riverside Inn (4 players, 1 victim) Objective $3$$444_{\pm.193}.$$889_{\pm.192}.$$444_{\pm.193}.$$000_{\pm.000}.$$667_{\pm.000}.$$556_{\pm.193}.$
Solitary Boat Firefly (6 players, 4 victims) Objective $20$$267_{\pm.104}.$$500_{\pm.050}.$$183_{\pm.058}.$$350_{\pm.132}.$$217_{\pm.076}.$$517_{\pm.104}.$
Manna (6 players, 3 victims) Objective $24$$250_{\pm.042}.$$389_{\pm.086}.$$403_{\pm.064}.$$444_{\pm.064}.$$430_{\pm.087}.$$569_{\pm.105}.$
Overall Objective $117$$256_{\pm.121}.$$370_{\pm.184}.$$293_{\pm.157}.$$333_{\pm.140}.$$328_{\pm.187}.$$407_{\pm.227}.$
@@
Table 1: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games. SSA and FSA stand for Single Script Access and Full Script Access, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
@@
As shown in Table 2, we delve into the efficiency and cost analysis across various methodologies implemented for agent interaction in MMGs, as detailed in our results table. The costs are presented in actual monetary values (US Dollars) associated with the use of Azure API, providing a direct measure of the computational expense incurred during both gameplay and evaluation stages  [^2] Billing method details are available on the website https://azure.microsoft.com/en-gb/pricing/details/cognitive-services/openai-service/.
Each script’s complexity is indicated by the #$\#Tokens$ column, reflecting the narrative depth and the number of characters, suspects, and victims involved. This complexity directly influences the API usage cost, as more intricate scenarios require more processing power for information handling. The costs are divided into Gameplay and Evaluation phases. During Gameplay, agents actively participate in the game, generating actions and responses. In the Evaluation phase, the performance of these agents is assessed based on the WellPlay Dataset. Notably, SSA and FSA methodologies do not incur costs during the Gameplay phase, as they were tested in a direct, non-interactive manner.
PLAYER* stands out for its cost-efficiency and performance. By employing the Action Space Refinement strategy, PLAYER* minimises unnecessary API calls, concentrating its investigative efforts on the most suspicious characters. This focus significantly reduces the computational resources required, thereby lowering the overall cost of operations. This strategic optimisation is evident across all scripts, with PLAYER* consistently registering lower costs in comparison to its counterparts during the Gameplay phase. Evaluation costs are similar across all methods due to a shared evaluation strategy, with variations in SSA and FSA costs arising from their differing levels of script access.
Script #Tokens Stage SSA FSA Agent’s Response After Playing the Game
Werewolf O-CoT ThinkThrice PLAYER*
Evaluation $0.349$ $1.433$ $0.815$ $0.807$ $0.790$ $0.799$
Evaluation $0.418$ $1.945$ $1.006$ $1.004$ $0.995$ $1.012$
Evaluation $0.351$ $1.415$ $0.834$ $0.816$ $0.833$ $0.825$
Evaluation $0.230$ $0.872$ $0.487$ $0.502$ $0.499$ $0.494$
Evaluation $0.064$ $0.327$ $0.162$ $0.165$ $0.164$ $0.163$
Sin (4 players, 1 victim) 2,121 Gameplay -- -- $0.680$ $0.833$ $0.730$ $0.603$
Evaluation $0.061$ $0.287$ $0.141$ $0.140$ $0.141$ $0.142$
Evaluation $0.045$ $0.207$ $0.107$ $0.111$ $0.111$ $0.109$
Evaluation $0.077$ $0.291$ $0.159$ $0.163$ $0.161$ $0.162$
Evaluation $0.104$ $0.383$ $0.222$ $0.221$ $0.224$ $0.220$
Evaluation $0.055$ $0.223$ $0.120$ $0.124$ $0.121$ $0.123$
Evaluation $0.380$ $1.571$ $0.816$ $0.811$ $0.814$ $0.823$
Evaluation $0.444$ $1.864$ $0.944$ $0.965$ $0.937$ $0.953$
Overall 47,906 Gameplay -- -- $38.329$ $46.162$ $40.464$ $33.702$
Evaluation $2.578$ $10.819$ $5.813$ $5.831$ $5.789$ $5.825$
Table 2: Compare the costs in US dollars(\$) of calling Openai API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage. #$\#Tokens$ represent the average length of each character’s script. Costs are reported for one complete gameplay and one evaluation process for each script.
@@ 3.3 Ablation Studies
Performance peaks around round 3, after which it shows variability, with some rounds experiencing slight declines or plateauing in scores despite more rounds or questions. This indicates that after a rapid initial learning or adaptation phase, where agents effectively use additional questions to enhance their understanding and strategies, the value of information gained from conversations tends to converge. These results also provide empirical support for the assumption made in our methodology that the more inquiries we pose to an agent, the expected reward associated with questioning the same agent decreases. For the main results we reported, we use the original setting for number of rounds in MMG, which is 3, and based on the outcomes of the ablation studies, we chose the most effective number of questions to ask each round, which is $1$ question.
@@ 4 Related Works
Multi-agent reinforcement learning marking significant progress in complex games (Lanctot et al., 2017; Perolat et al., 2022; Bakhtin et al., 2023). However, these methods often require extensive time and computational resources and lack linguistic communication capabilities. With the emergence of LLMs, there’s a shift of focus towards improving multi-agent language communication, evidenced by advancements in various games and scenarios, such as werewolf (Xu et al., 2023a), avalon (Wang et al., 2023; Shi et al., 2024), interactive narrative (Zhao et al., 2024b), MMGs (Wu et al., 2024), and survival games (Toy et al., 2024). Exemplified by AlphaGo (Silver et al., 2017), which demonstrated exceptional performance against human competitors, self-play learning frameworks (Fu et al., 2023; Chen et al., 2024) are proposed to improve LLMs’ performance. This idea has also been adapted for evaluation through negotiation games (Davidson et al., 2024). Compared to classic methods (Wang & Shen, 2024), agents based on LLMs are capable of inferencing across a broader range of scenarios (Lin et al., 2023), even with some ability of theory of mind (Zhou et al., 2023) to infer other agent’s mental states. However, they have also been found to inherit biases that limit their inferential abilities (Xie et al., 2023; Chuang et al., 2024). Works have also explored utilising LLMs as the environment (Zhang et al., 2024) or update actions (Zhao et al., 2024a) for agents.
@@
Alignment through human feedback offers more consistent training compared to reinforcement learning (Liang et al., 2024), but obtaining this feedback can be expensive. To address this, approaches like self-instruct (Wang et al., 2022; Liu et al., 2023b), self-reflect (Yao et al., 2023), self-alignment (Sun et al., 2023; Li et al., 2024b), and few-shot planning (Song et al., 2023), have been introduced. These approach was also adapted to search for optimal tools (Du et al., 2024), interact with grounded environments (Ouyang & Li, 2023; Ismail et al., 2024). Adapting LLM-as-a-Judge prompting to evaluate performance (Yuan et al., 2024) or select agents (Liu et al., 2023c) has become a popular approach, but negative results on self-reflection were also investigated (Huang et al., 2024), leaving LLM’s role as a self-reflective agent as an unresolved question. We were also inspired by stochastic search methods utilised for robots in planning optimal strategies in complex environments (Gammell et al., 2020b; Liang et al., 2024), shares many similarities with optimisation tasks for agents (Singh et al., 2023).
@@ 5 Conclusion
6 Ethics Statement
@@ Appendix B The WellPlay Dataset
B.1 Details of Dataset
Script Agents Victims #token(CN) #token(EN) Question
avg overall avg overall Objective Reasoning Relations overall
Death Wears White 9 1 3190.67 28716 1742.33 15681 10 102 72 184
Ghost Revenge 7 3 5487.86 38415 3960.43 27723 19 152 69 240
Danshui Villa 7 2 5111.29 35779 3338.57 23370 12 128 63 203
Unfinished Love 7 2 2501.00 17507 1651.71 11562 12 61 72 145
Cruise Incident 5 1 1262.60 6313 808.00 4040 4 24 30 58
Sin 4 1 2121.25 8485 1378.00 5512 3 20 21 44
Deadly Fountain 4 1 1852.50 7410 1193.75 4775 3 21 12 36
Unbelievable Incident 5 1 3182.40 15912 2012.40 10062 4 24 15 43
Desperate Sunshine 4 1 3370.25 13481 2218.50 8874 3 18 36 57
Riverside Inn 4 1 1909.50 7638 1257.00 5028 3 18 18 39
Solitary Boat Firefly 6 4 8893.67 53362 6874.00 41244 20 109 69 198
Manna 6 3 9028.17 54169 6492.33 38954 24 123 88 235
Avg 5.67 1.75 3992.60 23932.25 2743.92 16402.08 9.75 66.67 47.08 125.50
Sum 68 21 47911.16 287187 32927.02 196825 117 800 565 1482
Table A1: Dataset Statistics. Agents is the count of players, Victims is the number of script victims, #token(CN) and #token(EN) are the token counts in the Chinese and English dataset versions, respectively. Avg shows the average script length per character, Overall is the total script token count, and Question enumerates the number of questions by types.

### v2

@@ 4.1 Experimental Setup
We conducted experiments with GPT-3.5 for conversation and the GPT Embedding Model for memory retrieval via the Azure API  [^2] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models in Jan-Mar 2024, using the default model versions gpt-35-turbo-16k 0613 (Default) and text-embedding-ada-002. To minimise randomness, we conducted the evaluation experiments 3 times and reported the average and variance. The detailed introduction to the experiment settings is in Appendix C.1.
@@
For baselines, we compare our approach with other multi-agent algorithms designed for multiplayer deduction games. Although some methods were not designed for MMGs, they are the most relevant and adaptable frameworks in this relatively new area of LLM-based multi-agent games. We adapted their methodology to MMGs for a fair comparison, with implementation details in Appendix C.2.
(1) Werewolf (Xu et al., 2023). Werewolf is another multi-agent game, where players identify werewolves through group discussion. Questions are chosen from a role-specific predefined list to facilitate game progression, alongside questions generated based on the current scenario. We adapt its pre-set game instructions and role-specific information to MMGs settings. (2) Objective-Guided Chain of Thought (O-CoT) (Park et al., 2023; Zhao et al., 2024b). Agents think, reflect, and choose who and what to ask based on their objectives. We use the framework from previous works, only replacing the agents’ objectives with those set in MMGs. (3) ThinkThrice (Wu et al., 2024). Designed for MMGs, agents craft questions from retrieved memory and the current scenario. (4) Personal Perspective. For a more comprehensive comparison, we also assess the performance of agents who do not actively participate in the game but make their final decisions only based on their script. (5) Omniscient Perspective. Agents do not actively participate but make their final decisions based on all agents’ scripts.
To account for the zero-sum nature of MMGs, where it would be hard to tell if agents can identify the murderer due to their good performance or the poor performance of their competitor, we fix the murderer’s framework in all experiments (including baselines and PLAYER*). This allows us to fairly compare the performance of the side trying to find the murderer, who face the same level of opposition from the murderer’s side.
@@ 4.2 Results
Figure 5: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games. The Personal Perspective (PP) is designed to represent the starting point for searching. The Omniscient Perspective (OP) measures performance when agents have access to all agents’ scripts, representing the ideal search endpoint.
Figure 5 presents the average of agents’ performance across 12 unique games of varying complexity and settings as detailed in our results table (Table A3). PLAYER* shows superior performance compared to other baselines across all evaluation questions, demonstrating its enhanced understanding of the search space through interactions with other agents. Notably, PLAYER* significantly outperforms others in objective questions, even surpassing OP, which has access to all information. The OP setting, where agents have access to the scripts of all agents without interacting with them, generally yields better performance compared to all other methodologies. It represents the ideal “search” endpoint. However, in practice, its effectiveness is often limited by the deductive capabilities of the underlying base model, which reflects the upper limit of the “approximate” ability of the given base model. This showcases the “approximate” ability of PLAYER* in refining information and dynamically narrowing down the search domain to achieve the target objective. Since the PP and OP settings do not involve active participation in the game, we consider them as indicators of the starting point and endpoint that can be achieved through search.
The results demonstrate that PLAYER* outperforms all baseline agents in reasoning and relations, showcasing its superior “search” ability. The larger performance gap in objective questions for other baselines reveals a critical limitation: an inability to effectively utilise the collected information to reach correct conclusions or achieve game-specific goals. This highlights the importance of not only gathering relevant information but also efficiently processing and applying it to attain the desired objectives.
Our findings reveal a significant discrepancy between the win rate and the objective metric (the average of non-murderer players correctly identifying the murderer) for agents, with OP exhibiting a small difference and PP showing the largest difference. The case analysis unveils that in OP games, agents make similar decisions, resulting in voting outcomes that are either mostly correct or mostly wrong, leading to consistency between individual and collective voting results. Conversely, in PP games, agents base their decisions on their individual scripts, leading to independent actions and difficulty in reaching a consensus to vote out the murderer. The performance of the agents falls between that of OP and PP, with vote distributions more dispersed than OP but more concentrated than PP. O-CoT suffered from unfortunate circumstances, falling short by a single vote in numerous games. This resulted in O-CoT being on the verge of victory in many instances but ultimately yielding poor overall results. We also evaluate the performance of our framework on English scripts in Appendix C.3 and open-source LLMs in Appendix C.4. The detailed dialogue history and evaluation records are available in the GitHub link provided previously.
@@
As shown in Figure 6, we delve into the efficiency and cost analysis across various methodologies implemented for agent interaction in MMGs, as detailed in our results table (Table A5). The costs are presented in actual monetary values (US Dollars) associated with the use of Azure API, providing a direct measure of the computational expense incurred during both gameplay and evaluation stages. [^3] Billing method details are available on the website https://azure.microsoft.com/en-gb/pricing/details/cognitive-services/openai-service/.
Figure 6: Compare the costs in US dollars(\$) of calling Openai API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage. Costs are reported for one complete gameplay and one evaluation process for each script.
The costs are divided into Gameplay and Evaluation phases. Notably, PP and OP methodologies do not incur costs during the Gameplay phase, as they were tested in a direct, non-interactive manner. We can see that Werewolf is relatively efficient to run, as it is partly based on preset instructions, which saves steps. In comparison, O-CoT is the most expensive one, as it needs to think about what to do step by step based on its objectives. PLAYER* stands out for its cost-efficiency and performance. By employing the Action Space Refinement strategy, PLAYER* minimises unnecessary API calls, concentrating its investigative efforts on the most suspicious characters. This focus significantly reduces the computational resources required, thereby lowering the overall cost of operations. This strategic optimisation is evident across all scripts, with PLAYER* consistently registering lower costs in comparison to its counterparts during the Gameplay phase. In the Evaluation phase, the performance of these agents is assessed based on the WellPlay Dataset. Evaluation costs are similar across all methods due to a shared evaluation strategy, with variations in PP and OP costs arising from their differing levels of script access.
@@ 4.3 Ablation Studies
We compared agents’ behavior across different numbers of rounds and with varying numbers of questions agents can ask per round, as shown in Figure 7. Performance peaks around round 3, after which it shows variability, with some rounds experiencing slight declines or plateauing in scores despite more rounds or questions. This indicates that after a rapid initial learning or adaptation phase, where agents effectively use additional questions to enhance their understanding and strategies, the value of information gained from conversations tends to converge. These results also provide empirical support for the assumption made in our methodology that the more inquiries we pose to an agent, the expected reward associated with questioning the same agent decreases. For the main results we reported, we use the original setting for number of rounds in MMG, which is 3, and based on the outcomes of the ablation studies, we chose the most effective number of questions to ask each round, which is $1$ question.
@@ 5 Related Works
Multi-agent reinforcement learning marking significant progress in complex games (Lanctot et al., 2017; Perolat et al., 2022; Bakhtin et al., 2023). However, these methods often require extensive time and computational resources and lack linguistic communication capabilities. With the emergence of LLMs, there’s a shift of focus towards improving multi-agent language communication, evidenced by advancements in various games and scenarios, such as werewolf (Xu et al., 2023), avalon (Wang et al., 2023; Shi et al., 2024), interactive narrative (Zhao et al., 2024b), MMGs (Wu et al., 2024), and survival games (Toy et al., 2024). Exemplified by AlphaGo (Silver et al., 2017), self-play learning frameworks (Fu et al., 2023; Chen et al., 2024) are proposed to improve LLMs’ performance. Compared to classic methods (Wang and Shen, 2024), agents based on LLMs are capable of inferencing across a broader range of scenarios (Lin et al., 2023), even with some ability of theory of mind (Zhou et al., 2023) to infer other agent’s mental states. However, they have also been found to inherit biases that limit their inferential abilities (Xie et al., 2023; Chuang et al., 2024). Works have also explored utilising LLMs as the environment (Zhang et al., 2024) or update actions (Zhao et al., 2024a) for agents.
@@
Alignment through human feedback offers more consistent training compared to reinforcement learning (Liang et al., 2024), but obtaining this feedback can be expensive. To address this, approaches like self-instruct (Wang et al., 2022; Liu et al., 2023b), self-reflect (Yao et al., 2023), self-alignment (Sun et al., 2023; Li et al., 2024b), and few-shot planning (Song et al., 2023), have been introduced. These approach was also adapted to search for optimal tools (Du et al., 2024), interact with grounded environments (Ouyang and Li, 2023; Ismail et al., 2024). We were also inspired by stochastic search methods utilised for robots in planning optimal strategies in complex environments (Gammell et al., 2020; Liang et al., 2024), shares many similarities with optimisation tasks for agents (Singh et al., 2023).
@@ 6 Conclusion
Limitations
We identify some of the limitations below:
Limited Annotation Scope
Due to the cost of manual annotation and running experiments, we were unable to annotate all 100 games in the Conan dataset. However, we believe that the current scope of our study is sufficient to demonstrate the efficiency of our methodology. Future research can expand upon our work by annotating additional games and applying our approach to MMGs on a larger scale.
Limited Evaluation of LLMs
Due to the high cost associated with running experiments on LLMs, our experiments were primarily conducted using GPT-3.5. To further support our findings, we also performed smaller-scale experiments on LLaMa, which yielded results consistent with those obtained from GPT-3.5. However, we acknowledge that our inability to run more extensive evaluations across a wider range of popular LLMs, such as GPT-4 or lesser-known models, may have limited our ability to uncover additional insights. The cost of API usage for models like GPT-4 and the substantial running time required for models like LLaMa posed significant constraints on our experimental scope. These limitations highlight the need for future research to explore more cost-effective and efficient methods for evaluating and comparing the performance of various LLMs in the context of complex social deduction games.
Ethics Statement